OpenStudy (anonymous):

integrate (2x+1 )(x-1)/(x+3)

OpenStudy (anonymous):

\[f(x) = \frac{2x^{2}-x-1}{x+3}\]\[f'(x) = \frac{(x+3)(4x-1)-(1)(2x^{2}-x-1)}{(x+3)^{2}}\]\[f'(x) = \frac{(4x^{2}-x+12x-3)-(2x^{2}-x-1)}{(x+3)^{2}}\]\[f'(x) = \frac{2x^{2}+12x-2}{(x+3)^{2}}\]\[f'(x) = \frac{2x^{2}+12x-2}{(x+3)^{2}}\]

OpenStudy (anonymous):

i need help with integrating it not differentiating:)

OpenStudy (anonymous):

Oh sorry, I just realized I was differentiating it :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):



OpenStudy (anonymous):

how did u get the second part

OpenStudy (anonymous):

use the long division to divide (2x^2-x-1) over (x+3)

OpenStudy (anonymous):

oh lol thank you very much:)

OpenStudy (anonymous):

you are welcome.